A watch loses 2 minutes every hour. If it now shows 8:45, what time was it displaying exactly 30 hours earlier?
A) 3:15 B) 1:45 C) 2:15 D) 2:45 E) 3:45

Answer by josmiceli(19441)

1 hr before, it showed 7:47
2 hrs before, it showed 6:49
3 hrs before, it showed 5:51
. . . . .
30 hrs before, it showed in minutes 8:45 - 30*60 + 30x2 = 8:45 - 1800 + 60
8:45 - 1740 = 8:45 - 29 hrs
8:45 - 24 hrs = 8:45
8:45 - 5 hrs = 3:45
The answer is E)